Operation Summary for August 8, 2011 - August 10, 2011 |
Monday August 8 The day shift began with Operations monitoring the Tevatron (TEV) store 8961 (59.84E30) and the Recycler stash (74.5E10), with the Antiproton Source (Pbar) stacking (129.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-1010 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, MTA began taking beam around 10:45 AM. Operators reported at 11:40 AM that two Booster RF stations (BRF3 & 15) tripped off for the first of three times in 17 minutes; they reset. Operators reported at 12:20 PM that BRF15 & 16 were tripping off and being reset. Operators searched and secured the MI-31 enclosure at 12:53 PM. Operations terminated store 8961 at 1:43 PM, put the TEV into shot setup, and established store 8962 at 3:08 PM with an average initial luminosity (AIL) of 113.05E30. During the evening shift, after noticing a Booster alarm (B:HLRFRL), operators reported at 5:16 PM that they found Booster study events going out with no studiers around and the events were not placed properly in the Time Line Generator (TLG). At 5:38 PM, operators pulled the Booster study events out of the TLG. Operations disabled beam to T-1010 at 6:36 PM. Operations disabled beam to MTA at 6:48 PM. At 9:34 PM, the MI RF Watchdog alarm showed a 95 Degree LCW system pressure trip; it reset, but tripped off again. Operators investigated and found a loss of head pressure; operators changed out the nitrogen bottle that supplied the head pressure. A MI RF station (MIRF5) tripped off many times; operators bypassed the station. Operations terminated store 8962 around 10:55 PM and put the TEV into shot setup. At 11:21 PM, some Pelletron vacuum pumps (IPD03, IPA07, and IPA03) tripped off; they reset. Operators reported at 11:53 PM that the Pelletron suffered a full discharge while attempting to run beam; the vacuum was recovering slowly. Tuesday August 9 The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (21.2E10), with Pbar stacking (158.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 8963 at 12:23 AM with an AIL of 112.91E30. A Linac RF station (LRF4) tripped off at 12:41 AM; it reset. At 1:49 AM, the NuMI beam permit tripped off on an incomplete MIRF gap envelope; operators investigated, but found no RF stations off. At 3:42 AM, LRF1 tripped off and two minutes later LRF5 tripped off; they both reset. Operations established beam to T-1010 at 4:14 AM. MIRF11 tripped off; it reset. The Neutron Therapy Facility (NTF) began its startup at 7:39 AM to treat a patient later in the morning. During the day shift, Operations terminated store 8963 at 8:37 AM and put the TEV into shot setup. Water Works personnel reported at 9:05 AM that they were filling the TEV LCW system. At 9:21 AM, a Switchyard power supply (S:HP3DS) tripped off; operators notified EE Support personnel. Operators began preparing the Pelletron for an access at 9:57 AM. Operations established store 8964 at 9:58 AM with an AIL of 124.2E30. A Booster vacuum pump (B:IPRF17) tripped off and brought down BRF7 & 8; everything reset. At 11:17 AM, operators reported that a LCW valve (V131) at CUB developed a bad leak. Water Works personnel will take the columns offline and make repairs. At 11:36 AM, the NuMI beam permit tripped off for the first of four times in ten minutes due to a MI incomplete RFGE indication, but all the RF stations were on and everything reset. MI personnel began a coalescing study at noon. RF personnel continued to work on MIRF11. Water Works reported that the CUB valve repair was halted, the leak had stopped, and column II was brought online at 2:48 PM. Operators reported that BRF7’s power supply was tripping off; a Booster RF person investigated. During the evening shift, Operations disabled beam to T-1010 at 5:33 PM. Operations terminated store 8964 at 6:03 PM, put the TEV into shot setup, and established store 8965 with an AIL of 128.94E30. Operators reported at 11:02 PM that the resistivity of the Pbar 95 Degree LCW system was dropping; FESS personnel investigated. Wednesday August 10 The midnight shift began with Operations monitoring store 8965 (59.18E30) and the stash (158.4E10), with Pbar stacking (88.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators prepared the Meson M01 enclosure for access at 12:32 AM to investigate a sump pit high level alarm; FESS personnel reported that a faulty power breaker was the cause of the sump pump not working and would be replaced during the day shift. FESS personnel reported at 2:32 AM that they had replaces five deionizing bottles for the Pbar 95 Degree LCW system. Operations terminated store 8965 at 3:33 AM, put the TEV into shot setup, and established store 8966 at 4:50 AM with an AIL of 129.4E30. At 3:57 AM, BRF5 stopped putting out a gap envelope; operators bypassed the station and brought BRF19 on in its place. Operations established beam to T-1010 at 6:03 AM. Operators prepared MI-31 for access at 6:37 AM; Pelletron personnel reported that they were continuing their repairs. The Plans for Wednesday The plans for today are to stack, stash, and store. Pelletron repairs continue. The Run Coordinator will allow CUB personnel to reconfigure the columns back to their normal running conditions, for EE Support personnel to replace the PLC for the Switchyard power supply HPS3DS, and for Linac, Booster, and MI studies. |
|
More Information Comments and Suggestions |
| last modified 08/10/2011 email Fermilab |
| Security, Privacy, Legal | |